OpenStudy (anonymous):

@whpalmer4 Use log\[\log_{2} 3 \approx 1.5850 and \log_{2} 5\approx 2.3219 \to approximate the value of each expression.\]

OpenStudy (anonymous):

1) \[\log_{2} 25\]

OpenStudy (whpalmer4):

how could you write 25 using 5 and 3 and multiplication and division?

OpenStudy (anonymous):

well is actually suppose to represent two squiggly lines sort of like an equal sign

OpenStudy (whpalmer4):

Yeah, \(\approx\) means approximately But the key to this problem is figuring out how to express 25 using only 5 and 3 and multiplication and division (may not have to use all of them)

OpenStudy (whpalmer4):

Hint: what is 5*5?

OpenStudy (anonymous):

25

OpenStudy (preetha):

Whpalmer - good work there!

OpenStudy (whpalmer4):

Right. Remember the property of logarithms involving multiplication? \[\log (a*b) = \log a + \log b\]

OpenStudy (whpalmer4):

@Preetha thanks, boss ;-)

OpenStudy (anonymous):

but that's not the answer right

OpenStudy (whpalmer4):

We can write \[\log 25 = \log (5*5) = \log 5 + \log 5\]right? Works for any base...

OpenStudy (anonymous):

but only for this particular problem

OpenStudy (whpalmer4):

No, it's always true. For logarithms of any base \(b\), \[\log_b (x*x) = \log_b x +\log_b x = 2\log_b x\]

OpenStudy (anonymous):

no no I understand that rule but as for the 25 and 5

OpenStudy (whpalmer4):

we happen to know the value of \(\log_2 5\), so we can use that to our advantage: \[\log_2 25 = \log_2 (5*5) = \log_2 5 + \log_2 5 = 2*\log_2 5 \approx 2*2.2319 \approx 4.6439\]

OpenStudy (whpalmer4):

We could also write it like this: \[\log_2 25 = \log_2 5^2 = 2\log_2 5\]because \[\log_b x^a = a\log_b x\]

OpenStudy (anonymous):

ooohhhh ok I get it now

OpenStudy (whpalmer4):

really, the whole trick here was realizing how to "construct" 25 from the building blocks we have, 5 and 3

OpenStudy (whpalmer4):

\[\log_2 75 = \log_2 (3*25) = \log_2 3 + 2\log_2 5 = 1.5850 + 4.6439 = 6.22882\] (for example)

OpenStudy (anonymous):

and since 5 is a factor of 25 we choose that as the number they gave us

OpenStudy (anonymous):

I got cha :)

OpenStudy (whpalmer4):

as I mentioned earlier, I memorized a handful of log values, and I can use them like tinker toys to estimate logs of other numbers, just like I did with 25 and 75 here. there's another property of logs which is useful, called the base change property: \[\log_b x = \frac{\log_a x}{\log_a b}\] So if you know the log in one base, you can find it in another base by some simple arithmetic.

OpenStudy (whpalmer4):

For example, if you wanted to compute the log base 2 of 25, but didn't have those values handy, but did have a calculator that does \(\log_{10}\), you could do this: \[\log_2 25 = \frac{\log_{10} 25}{\log_{10}2} = \frac{1.39794}{0.30103} = \](go ahead, do it!)

OpenStudy (anonymous):

i think she was telling us about that

OpenStudy (whpalmer4):

so, what do you get when you do that division on your calculator?

OpenStudy (anonymous):

4.6439

OpenStudy (whpalmer4):

Right. Does that answer look familiar? :-)

OpenStudy (anonymous):

yep lol we just took a short cut :)

OpenStudy (whpalmer4):

The fun thing about problems with logarithms (to the extent that there is a fun thing) is that often there are many different ways to get to the answer. It can also be a frustrating thing, if you can't figure out how to do any of them!

OpenStudy (anonymous):

exactly which was how I was before you taught me

OpenStudy (whpalmer4):

Sounds like a satisfied customer :-)

OpenStudy (anonymous):

definitely

OpenStudy (whpalmer4):

was that all of these problems, or have you done the rest of them while I've been faffing about here trying to juggle multiple questions?

OpenStudy (anonymous):

Oh trust me I have plenty more to do she gave us like 3 worksheet but I don't want to bother you anymore because I don't want you to get tired and other ppl may need help :)

OpenStudy (anonymous):

@whpalmer4

OpenStudy (whpalmer4):

Well, I keep hopping off to help others as they summon me (just like you do). It's more likely at this hour of the day (7PM my time) that you'll get tired than that I will :-)

OpenStudy (anonymous):

lol so ur saying I can ask you more lol

OpenStudy (anonymous):

@whpalmer4

OpenStudy (whpalmer4):

Yes. By the way, I get a notification each time you post in the thread, no need to also tag me unless you grow impatient :-)

OpenStudy (anonymous):

oh my bad lol I'm sorry ok here we go solve each inequality or equation. Round to the nearest ten-thousandth 9) 3^x>243